OpenStudy (anonymous):

Simplify. Assume that no denominator is 0 and that 0^0 is not considered. (4/3x)^3

OpenStudy (lgbasallote):

\[(\frac{a}{b})^c = \frac{a^c}{b^c}\] does that help?

OpenStudy (anonymous):

no i am totally confused with this can you please show me the steps so i can follow for others

OpenStudy (lgbasallote):

it just means if you have (4/3x)^3 then just cube the numerator and cube the denominator \[\large \frac{4^3}{(3x)^3}\] now can you do it?

OpenStudy (anonymous):

so then all i do is divide then ^ ?

OpenStudy (lgbasallote):

no...you cube

OpenStudy (lgbasallote):

4^3 = 4 x 4 x 4 (3x)^3 = 3x * 3x * 3x

OpenStudy (anonymous):

thats the answer ? i messed it, now i can see what it means

OpenStudy (lgbasallote):

you have to simplify it further

OpenStudy (lgbasallote):

what is 4 x 4 x 4?

OpenStudy (anonymous):

64

OpenStudy (lgbasallote):

and what is 3x * 3x * 3x?

OpenStudy (anonymous):

27

OpenStudy (lgbasallote):

you forgot x* x* x

OpenStudy (lgbasallote):

27 is just 3 x 3 x 3

OpenStudy (anonymous):

x^3

OpenStudy (lgbasallote):

therefore final answer is 64/27x^3

OpenStudy (anonymous):

oh my i understand it I feel stupid, it is so easy to understand. thank you

OpenStudy (lgbasallote):

no probs ^_^